OpenStudy (anonymous):

1. y - 2x - 6 =0 and y/2 - 3 = x 2. 3x + y = 0 and 2y - 4x = 10 3. 10y + 35 = 15x and 4 = 3x - 2y

OpenStudy (anonymous):

okay..so whad do i do? solve, graph....

OpenStudy (anonymous):

For the systems of linear equations in questions 1-3: Graph the two lines, and list the x-intercepts, y-intercepts, and points of intersection. I'll assess the answers as follows: 2 point of intersection; that is, the solution[s] [if any] 2 one line's x and y intercepts. 2 other line's x and y intercepts 2 correct graph In order to earn credit for the slope and intercepts, you must explicitly write each of these sentences, filling in the blanks with the correct values for that problem: A. For equation ___________ 1. The x-intercept is ___. 2. The y-intercept is ___. 3. The slope is ___. B. For equation ___________ 1. The x-intercept is ___. 2. The y-intercept is ___. 3. The slope is ___.

OpenStudy (anonymous):

wait..so i all ineed to do is find the x and y int of each? or graph also?

OpenStudy (anonymous):

yes and slope.

OpenStudy (anonymous):

equation a: xint: 3 yint:-6 slope:2

OpenStudy (anonymous):

do u want me to graph it?

OpenStudy (anonymous):

yea, but isn't there more to it? than just 1 answer?

OpenStudy (anonymous):

yea..tas only a

OpenStudy (anonymous):

do u want me to the others then graph?

OpenStudy (anonymous):

yes please?

OpenStudy (anonymous):

kay..lemme try..i hope iam not wrong...

OpenStudy (anonymous):

dunt really understand b: i got like (0,0) but dunno if its correct...

OpenStudy (anonymous):

oh boy lol

OpenStudy (anonymous):

srry..looks like i jinxed my self..lemme do the others..

OpenStudy (anonymous):

2a: slope=-3 no x or y int..

OpenStudy (anonymous):

2b: slope=2 xint= 5/2 yint:5

OpenStudy (anonymous):

3a: slope=1.5 xint= 7/3 yint: -3.5

OpenStudy (anonymous):

wait i dont understand the numbering 1. of both problems is? 2. 3.

OpenStudy (anonymous):

3b: slope=3/2 xint=4/3 yint=2

OpenStudy (anonymous):

your confusing me lol

OpenStudy (anonymous):

okay, so tare are 3 problems with 2 equatiosns, so the first equation is 1a the second one is 1b, then its 2a, 2b, 3a, 3b..do u understand

OpenStudy (anonymous):

lol ew dont get so frustrated. yes i do now

OpenStudy (anonymous):

LOL..im not frustrated..

OpenStudy (anonymous):

srry..but i cant graph them rite now...i gots to go....